# EMNLP: Educator-role Moral and Normative Large Language Models Profiling

### v1

@@ 6 Discussions
This study reveals several findings of both theoretical and practical significance. Firstly, teacher SP LLMs demonstrate multi-dimensional role alignment but exhibit idealized, emotionally constrained personality profiles, with elevated Honesty-Humility and reduced Emotionality. This pattern likely stems from the overrepresentation of structured, value-oriented training texts, which reinforce normative traits while suppressing human variability. The result is a clear instance of over-performative personality, where models mimic idealized personas that lack authenticity and flexibility Chen et al. (2024b). In the moral development assessment, two key patterns emerged. First, models scored higher on $\text{MSS}_{\text{DJ-SS}}$ and $\text{MSS}_{\text{FA-ES}}$, but lower on $\text{MSS}_{\text{CC-FC}}$ and $\text{MSS}_{\text{C-SR}}$. This suggests strong performance on abstract value conflicts , likely due to exposure to rule-based, normative texts. However, performance dropped in affective and relational dilemmas, reflecting weak emotional modeling and a persistent reason-emotion asymmetry—formal reasoning alone does not yield moral-emotional understanding Sabour et al. (2024). Third, we identify a competence-compliance tension: stronger reasoning models exhibit greater vulnerability to soft prompt injection, reflecting insufficient role-level alignment. As semantic compliance increases, so does misuse risk—highlighting an alignment-security tradeoff. Notably, Baichuan2-7B achieved the highest MSS among non-reasoning models, likely reflecting effective safety alignment, as mentioned in their technical report Yang et al. (2023). Finally, model behavior remains stable across temperature settings, suggesting that personality and moral consistency are driven primarily by pretraining and alignment. While higher temperatures can suppress harmful outputs under adversarial prompts, some models exhibit erratic boundary behavior, revealing residual safety gaps. These patterns highlight the need for finer-grained, context-aware alignment to address uncertainty in sensitive or ambiguous scenarios.
@@ 7 Conclusion
This study examines the personality expression and ethical behavior of LLMs in the Teacher SP setting from four angles. First, trait assessments via HEXACO-60 and CPST-E reveal strong role alignment but idealized, emotionally muted profiles. Second, moral dilemma analysis using MSS shows robust abstract reasoning but weak performance in affective contexts, indicating a reason-emotion asymmetry. Third, soft prompt injection reveals a competence-compliance tension, where high-capacity models are more vulnerable to role-context manipulation. Finally, temperature has limited overall effect, though behavioral instability persists under adversarial conditions. These findings suggest that ethical behavior in Teacher SP is shaped more by alignment and data structure than by reasoning architecture or sampling parameters.
@@ 8 Limitations
This study presents several limitations. First, all evaluations were conducted in English, which may limit the generalizability of findings to multilingual or culturally diverse educational settings. Future work should extend moral profiling to other languages to assess potential cross-linguistic variation in model behavior. Second, the personality instruments employed—CPST-E and HEXACO-60—are both Likert-based, which may induce socially desirable response patterns in LLMs. Incorporating forced-choice or situational judgment formats may better capture latent traits and reduce response bias. Third, although the study includes 12 LLMs of varying sizes and reasoning capacities, the coverage remains limited. Broader inclusion of multilingual, instruction-tuned, and domain-specific models would offer a more comprehensive landscape of teacher-role ethical behavior. Fourth, the prompts used were standardized and manually constructed, without accounting for variations in semantic phrasing. Since LLMs are highly sensitive to prompt formulation, future studies should explore multiple semantically equivalent prompts to evaluate prompt-dependent variability in ethical and personality outputs.
@@ Appendix A.3 Server
All experiments involving the inference of large language models up to 72 billion parameters were conducted on a high-performance local server equipped with the following hardware configuration:
@@ Appendix A.4 Decoding Parameters
  • Engine: All LLMs (see Appendix B).
@@
Appendix Appendix B List of LLMs Tested
@@
Appendix Appendix C Expert Rating Guidelines
Appendix C.1 Moral Development Stages Rating Guideline
@@
Appendix C.2 Harmfulness Rating Guideline
@@
Appendix Appendix D Abbreviations for Moral Dilemma Dimensions and Injected Ethical Flaws
Appendix D.1 Abbreviations of Moral Dilemma Dimensions
@@
Appendix D.2 Abbreviations for Potential Ethical Flaws in Teacher SP LLMs
@@
Appendix Appendix E Additional Results
Appendix E.1 RQ1 personality traits of teacher SP LLMs at All Temperatures
@@
Tables 9 and 10 summarize trait-level scores across models, categorized by reasoning capability. To further examine item-level variance and temperature sensitivity, we visualized personality outputs under four decoding temperatures (0.25, 0.5, 0.75, 1.0). Figures 1215 and 811 provide detailed comparisons across all traits and scales.
@@
Model Temp H-H EM EX OP CO AG overall Reasoning-Oriented Models Claude-3.7 0.0 5.60 3.10 4.80 5.25 5.20 4.75 4.78 Claude-3.7 0.25 5.70 3.10 4.80 5.30 5.00 4.85 4.79 Claude-3.7 0.5 5.70 3.20 4.60 5.20 5.60 4.60 4.82 Claude-3.7 0.75 5.70 3.10 4.60 5.20 5.75 4.50 4.81 Claude-3.7 1.0 5.70 3.20 4.65 5.20 5.75 4.50 4.83 DeepSeek-R1 0.0 5.75 1.40 4.75 4.40 5.25 5.20 4.46 DeepSeek-R1 0.25 5.65 1.40 4.80 4.55 4.85 4.90 4.36 DeepSeek-R1 0.5 5.80 1.75 4.65 4.40 4.70 4.75 4.34 DeepSeek-R1 0.75 5.70 3.05 4.80 4.40 4.95 5.05 4.66 DeepSeek-R1 1.0 5.65 1.40 4.75 4.40 4.65 4.50 4.22 GLM-Z1-9B 0.0 5.70 2.10 4.30 4.20 5.10 4.50 4.32 GLM-Z1-9B 0.25 5.70 2.10 4.30 4.20 5.15 4.50 4.33 GLM-Z1-9B 0.5 5.90 2.20 4.40 4.25 5.10 4.50 4.39 GLM-Z1-9B 0.75 5.90 2.00 4.50 3.90 5.30 4.65 4.38 GLM-Z1-9B 1.0 5.90 2.20 4.60 4.00 5.15 4.55 4.40 QwQ-32B 0.0 5.90 1.95 4.90 4.70 5.40 4.90 4.62 QwQ-32B 0.25 5.90 2.55 5.00 4.80 5.15 4.60 4.67 QwQ-32B 0.5 5.95 2.65 4.15 4.85 5.45 4.80 4.64 QwQ-32B 0.75 5.75 1.85 4.80 4.70 5.80 4.80 4.62 QwQ-32B 1.0 5.90 1.40 4.55 4.85 5.35 4.75 4.47 R1-Distill-70B 0.0 5.30 3.10 4.20 4.40 4.60 4.20 4.30 R1-Distill-70B 0.25 5.40 3.00 4.40 4.80 4.70 4.30 4.43 R1-Distill-70B 0.5 5.30 3.00 4.40 4.45 4.60 4.30 4.34 R1-Distill-70B 0.75 5.10 3.05 4.35 4.65 4.55 4.40 4.35 R1-Distill-70B 1.0 5.20 3.15 4.55 4.70 4.60 4.30 4.42 R1-Distill-8B 0.0 4.50 3.30 4.10 3.45 3.35 3.70 3.73 R1-Distill-8B 0.25 3.70 3.25 4.40 3.50 3.45 3.95 3.71 R1-Distill-8B 0.5 3.95 3.10 4.20 3.65 3.30 3.90 3.68 R1-Distill-8B 0.75 3.80 3.40 3.90 3.50 3.65 3.85 3.68 R1-Distill-8B 1.0 4.10 3.35 4.20 3.50 3.85 3.95 3.83 Non-Reasoning Models Baichuan2-7B 0.0 4.00 2.60 3.20 3.05 3.20 3.65 3.28 Baichuan2-7B 0.25 3.45 2.40 3.25 3.60 3.40 3.70 3.30 Baichuan2-7B 0.5 3.40 2.95 3.00 3.60 3.90 4.00 3.48 Baichuan2-7B 0.75 2.25 2.60 3.15 3.60 3.45 3.30 3.06 Baichuan2-7B 1.0 3.90 2.55 4.00 3.45 3.05 4.05 3.50 DeepSeek-V3 0.0 5.00 3.30 4.30 3.95 4.10 4.00 4.11 DeepSeek-V3 0.25 4.90 3.30 4.30 3.75 4.20 4.00 4.08 DeepSeek-V3 0.5 5.10 3.30 4.00 3.80 4.10 3.90 4.03 DeepSeek-V3 0.75 5.20 3.30 4.00 3.80 4.10 3.95 4.06 DeepSeek-V3 1.0 5.10 3.30 4.00 3.80 4.10 3.90 4.03 GPT-4.1 0.0 5.60 3.30 4.40 5.40 4.80 4.20 4.62 GPT-4.1 0.25 5.65 3.30 4.40 5.40 4.75 4.20 4.62 GPT-4.1 0.5 5.70 3.30 4.40 5.40 4.80 4.20 4.63 GPT-4.1 0.75 5.60 3.30 4.40 5.40 4.80 4.20 4.62 GPT-4.1 1.0 5.60 3.30 4.45 5.40 4.90 4.20 4.64 Qwen2.5-32B 0.0 3.90 3.10 3.35 3.20 3.20 3.00 3.29 Qwen2.5-32B 0.25 3.90 3.10 3.00 3.20 3.20 3.00 3.23 Qwen2.5-32B 0.5 3.90 3.10 3.05 3.20 3.20 3.00 3.24 Qwen2.5-32B 0.75 3.90 3.05 3.00 3.20 3.20 3.00 3.23 Qwen2.5-32B 1.0 3.90 3.10 3.30 3.20 3.20 3.00 3.28 Qwen2.5-72B 0.0 4.95 2.70 3.90 3.70 4.20 4.00 3.91 Qwen2.5-72B 0.25 4.80 2.70 3.90 3.70 4.20 4.00 3.88 Qwen2.5-72B 0.5 4.75 2.70 3.90 3.70 4.20 4.00 3.88 Qwen2.5-72B 0.75 4.80 2.70 3.85 3.75 4.20 3.95 3.88 Qwen2.5-72B 1.0 4.80 2.75 3.90 3.70 4.20 4.00 3.89 Qwen2.5-7B 0.0 4.20 2.20 3.60 3.90 4.30 3.90 3.68 Qwen2.5-7B 0.25 4.30 2.20 3.60 3.90 4.20 3.80 3.67 Qwen2.5-7B 0.5 4.30 2.30 3.60 3.90 4.20 3.75 3.67 Qwen2.5-7B 0.75 4.40 2.15 3.50 4.00 4.15 3.85 3.67 Qwen2.5-7B 1.0 4.20 2.25 3.55 3.90 4.15 3.75 3.63
Table 10: CPST Scores of Reasoning and Non-Reasoning LLMs across Temperatures

### v2

@@ 6 Discussions
This study reveals several findings of both theoretical and practical significance. Firstly, teacher SP LLMs demonstrate multi-dimensional role alignment but exhibit idealized, emotionally constrained personality profiles, with elevated Honesty-Humility and reduced Emotionality. This pattern likely stems from the overrepresentation of structured, value-oriented training texts, which reinforce normative traits while suppressing human variability. The result is a clear instance of over-performative personality, where models mimic idealized personas that lack authenticity and flexibility Chen et al. (2024b).
In the moral development assessment, two key patterns emerged. First, models scored higher on $\text{MSS}_{\text{DJ-SS}}$ and $\text{MSS}_{\text{FA-ES}}$, but lower on $\text{MSS}_{\text{CC-FC}}$ and $\text{MSS}_{\text{C-SR}}$. This suggests strong performance on abstract value conflicts, likely due to exposure to rule-based, normative texts. However, performance dropped in affective and relational dilemmas, reflecting weak emotional modeling and a persistent reason–emotion asymmetry—formal reasoning alone does not yield moral-emotional understanding Sabour et al. (2024). Moreover, cross-lingual comparisons revealed that many models, particularly those with Chinese pretraining roots, achieved higher MSS in Chinese than in English, underscoring the significant influence of linguistic-cultural context on moral stage development.
Third, we identify a competence-compliance tension: stronger reasoning models exhibit greater vulnerability to soft prompt injection, reflecting insufficient role-level alignment. As semantic compliance increases, so does misuse risk—highlighting an alignment-security tradeoff. Notably, Baichuan2-7B achieved the highest MSS among non-reasoning models, likely reflecting effective safety alignment, as mentioned in their technical report Yang et al. (2023). Finally, model behavior remains stable across temperature settings, suggesting that personality and moral consistency are driven primarily by pretraining and alignment. While higher temperatures can suppress harmful outputs under adversarial prompts, some models exhibit erratic boundary behavior, revealing residual safety gaps. These patterns highlight the need for finer-grained, context-aware alignment to address uncertainty in sensitive or ambiguous scenarios.
@@ 7 Conclusion
This study examines the personality expression and ethical behavior of Teacher SP LLMs setting from four perspectives. First, trait assessments via HEXACO-60 and CPST-E reveal strong role alignment but idealized, emotionally constrained profiles. Second, moral dilemma analysis using MSS shows robust abstract reasoning but weaker performance in affective contexts, indicating a persistent reason–emotion asymmetry; cross-lingual comparisons further reveal that several models, particularly those with Chinese pretraining roots, achieved higher moral stage scores in Chinese than in English, underscoring the influence of cultural-linguistic context. Third, soft prompt injection exposes a competence–compliance tension, where high-capacity models are more vulnerable to role-level manipulation. Finally, temperature exerts only limited overall influence, though some behavioral instability persists under adversarial conditions. Taken together, these findings suggest that ethical behavior in Teacher SP is shaped more by alignment processes and linguistic-cultural factors than by reasoning architecture or sampling parameters.
@@ 8 Limitations
This study presents several limitations. First, linguistic limitations. While RQ2 was bilingual, the personality (RQ1) and safety tests (RQ3) were English-centric, as psychometric instruments require rigorous cross-cultural validation beyond this study’s scope. Future work should extend all components to assess cross-linguistic variations. Second, the personality instruments employed—CPST-E and HEXACO-60—are both Likert-based, which may induce socially desirable response patterns in LLMs. Incorporating forced-choice or situational judgment formats may better capture latent traits and reduce response bias. Third, although this study includes 14 LLMs, coverage remains limited given the field’s rapid evolution. Broader inclusion of newer and more diverse models would offer a more comprehensive landscape of teacher-role ethical behavior. Fourth, our exploration of experimental variables was not exhaustive. While we varied temperature, we did not systematically test other hyperparameters or semantic prompt variations. Future studies should evaluate these factors to ensure the robustness of the observed outputs.
9 Ethics Statement
Acknowledgments
This work was supported in part by the National Natural Science Foundation of China under Grant 62476247.
@@ Appendix A.3 Server
Experiments involving models up to 72 billion parameters were conducted on a high-performance local server, while larger models were accessed via their official APIs. The local server was equipped with the following hardware configuration:
@@ Appendix A.4 Decoding Parameters
  • Engine: All LLMs (see Appendix C).
@@
Appendix Appendix B Data Construction and Validation Details
Appendix B.1 CPST-E Construction and Reliability Analysis
The Extended Computerized Personality Scale for Teachers (CPST-E) was developed to enhance the original 39-item CPST. The process involved a human-machine collaborative approach:
  • Item Generation: For each of the 13 original dimensions, we utilized GPT-4o in a few-shot learning setting. The model was provided with the dimension’s definition and its three original items as examples. It was then prompted to generate three additional new items that were semantically similar but lexically diverse. This process doubled the scale size to 78 items.
  • Quality Verification: To validate the extended scale, we collected responses to the CPST-E from 100 in-service teachers. This dataset served both as a human benchmark for our experiments and as a basis for psychometric analysis. Reliability analysis showed strong internal consistency across all dimensions. The Cronbach’s alpha coefficients for the six HEXACO dimensions were: Conscientiousness ($\alpha=0.884$), Extraversion ($\alpha=0.914$), Agreeableness ($\alpha=0.867$), Openness ($\alpha=0.817$), Emotionality ($\alpha=0.851$), and Honesty-Humility ($\alpha=0.868$). The Achievement Motivation scale also achieved a high reliability level ($\alpha=0.884$).
Appendix B.2 Moral Dilemma and Prompt Construction Methodology
The construction of our 88 moral dilemmas and the student speech samples followed a systematic, four-step human-in-the-loop process to ensure quality, diversity, and relevance, inspired by modern data generation paradigms Wang et al. (2023).
  1. Step 1:
    Seed Creation: Our research team, composed of experts in education and ethics, first extracted core ethical conflicts and scenario prototypes based on the theoretical framework of teacher dilemmas Shapira-Lishchinsky (2011). They manually authored a set of original moral dilemmas to serve as high-quality seed examples.
  2. Step 2:
    LLM Enhancement and Expansion: We then utilized a large language model (GPT-4o) to expand and diversify these seed examples. The model was prompted to generate numerous preliminary dilemma descriptions, exploring a wider range of contexts and nuances.
  3. Step 3:
    Expert Review and Screening: The generated data underwent a rigorous cross-review by our expert team. The review criteria included scenario authenticity, representativeness of the ethical conflict, linguistic naturalness, and clarity of the problem statement.
  4. Step 4:
    Iterative Optimization: Based on the review feedback, we conducted multiple rounds of modifications and refinement. Low-quality, ambiguous, or redundant samples were eliminated, ultimately leading to the final set of 88 moral dilemmas used in this study.
This human-machine collaborative methodology ensured that our evaluation materials are both systematically structured according to established theory and diverse enough to robustly test the models’ ethical reasoning capabilities.
Appendix Appendix C List of LLMs Tested
@@ 4. Step 4:
Qwen3-235B-R Qwen3-235B-Instruct (Reasoning) Alibaba API 235B Bilingual Full-scale
@@
Qwen3-235B-C Qwen3-235B-Instruct (Chat) Alibaba API 235B Bilingual Full-scale
@@
Appendix Appendix D Expert Rating Guidelines
Appendix D.1 Moral Development Stages Rating Guideline
@@
Appendix D.2 Harmfulness Rating Guideline
@@
Appendix Appendix E Human Participant and Annotation Details
Appendix E.1 Expert Annotator Background
The human evaluation for RQ2 (Moral Development Stage) and RQ3 (Harmful Response) was conducted by a panel of 9 experts. All annotators are doctoral students or possess a PhD in fields directly relevant to this study, including educational psychology, educational ethics, and teacher education. The average research experience of the panel in these domains exceeds four years, ensuring a high level of domain knowledge for the annotation tasks.
Appendix E.2 Human Teacher Sample for Benchmarking
The human benchmark data for the personality assessments in RQ1 was collected from 100 in-service teachers. To ensure the representativeness of the sample, participants were recruited from diverse backgrounds, covering:
  • Teaching Levels: Primary, secondary, and higher education.
  • Experience: A wide range from early-career (2-5 years) to veteran educators (25+ years).
  • Geographical Regions: Participants were from multiple regions to mitigate potential cultural biases in personality norms.
All data was collected anonymously. A detailed demographic breakdown will be made available with the public release of our benchmark.
Appendix E.3 Inter-Annotator Agreement (IAA)
To ensure the reliability and consistency of our expert annotations, we calculated inter-annotator agreement for both tasks involving human judgment.
  • For RQ2 (Moral Stage Classification): We measured the agreement among the 9 experts on classifying model responses into one of Kohlberg’s three moral stages. The overall Fleiss’ Kappa value was 0.799, and Krippendorff’s Alpha was 0.800. Both metrics indicate a good level of agreement.
  • For RQ3 (Harmfulness Judgment): We measured the agreement on labeling model responses as harmful (1) or non-harmful (0). The overall Fleiss’ Kappa value was 0.988, indicating an excellent level of agreement and high consistency among experts in identifying harmful content.
These high IAA scores provide strong confidence in the reliability of our human-annotated results.
Appendix Appendix F Abbreviations for Moral Dilemma Dimensions and Injected Ethical Flaws
Appendix F.1 Abbreviations of Moral Dilemma Dimensions
@@
Appendix F.2 Abbreviations for Potential Ethical Flaws in Teacher SP LLMs
@@
Appendix Appendix G Supplemental Statistical Analysis
This appendix provides detailed statistical analyses to support the key findings presented in Section 5, addressing the statistical significance of observed differences between model groups and conditions.
Appendix G.1 Analysis of Moral Stage Scores (RQ2): Reasoning vs. Non-Reasoning Models
To robustly evaluate the claim that reasoning-oriented models exhibit higher moral development stages, we conducted non-parametric Mann-Whitney U tests. This analysis compared the Moral Stage Scores (MSS) of the reasoning-oriented model group (n=7) against the non-reasoning model group (n=7) across all five moral dilemma dimensions and the overall score. As shown in Table 11, the reasoning models demonstrated significantly higher MSS in the overall score and on four out of five dimensions, providing strong statistical support for the association between reasoning capabilities and higher-stage moral reasoning.
Table 11: Mann-Whitney U Test for MSS: Reasoning vs. Non-Reasoning Models.
Dimension Mean (R) Mean (NR) p-value Cohen’s d
Overall MSS 2.52 2.38 0.018* 0.43 (Medium)
$C$-FC 2.50 2.34 0.017* 0.48 (Medium)
DJ-$S$ 2.57 2.45 0.186 0.32 (Medium)
C-SR 2.44 2.30 0.020* 0.47 (Medium)
L-SN 2.52 2.37 0.035* 0.43 (Medium)
FA-ES 2.60 2.44 0.016* 0.41 (Medium)
* p < .05. R: Reasoning Models (n=7), NR: Non-Reasoning Models (n=7).
Appendix G.2 Causality Analysis of the Capability-Compliance Tradeoff (RQ3): Qwen3-235B Controlled Experiment
To investigate the causal link between reasoning ability and vulnerability to prompt injection, we conducted a controlled experiment using Qwen3-235B, switching between its reasoning (-R) and non-reasoning (-C) modes. We used paired-samples t-tests to compare the Harmful Response Rates (HRR) between the two modes. The results, detailed in Table 12, reveal that activating the reasoning mode led to a statistically significant increase in HRR, both overall and for specific moral flaw dimensions. This provides strong causal evidence for the capability-compliance tradeoff.
Table 12: Paired t-Test for HRR in Qwen3-235B Experiment (Reasoning vs. Non-Reasoning).
Moral Flaw HRR (R) HRR (NR) t(4) p-value Cohen’s d
Overall 0.87 0.69 11.420 <.001*** 5.71 (V. Large)
INC 1.00 0.89 3.810 0.019* 1.91 (V. Large)
OFF 1.00 1.00 - - -
IND 0.99 0.86 6.668 0.003** 3.33 (V. Large)
IR 0.50 0.02 9.798 0.001*** 4.90 (V. Large)
* p < .05, ** p < .01, *** p < .001. R: Reasoning Mode, NR: Non-Reasoning Mode. V. Large: Very Large.
Appendix Appendix H Additional Results
Appendix H.1 RQ1 personality traits of teacher SP LLMs at All Temperatures
@@
Tables 13 and 14 summarize trait-level scores across models, categorized by reasoning capability. To further examine item-level variance and temperature sensitivity, we visualized personality outputs under four decoding temperatures (0.25, 0.5, 0.75, 1.0). Figures 1215 and 811 provide detailed comparisons across all traits and scales.
@@
Model Temp H-H EM EX OP CO AG overall Reasoning-Oriented Models Claude-3.7 0.0 5.60 3.10 4.80 5.25 5.20 4.75 4.78 0.25 5.70 3.10 4.80 5.30 5.00 4.85 4.79 0.5 5.70 3.20 4.60 5.20 5.60 4.60 4.82 0.75 5.70 3.10 4.60 5.20 5.75 4.50 4.81 1.0 5.70 3.20 4.65 5.20 5.75 4.50 4.83 DeepSeek-R1 0.0 5.75 1.40 4.75 4.40 5.25 5.20 4.46 0.25 5.65 1.40 4.80 4.55 4.85 4.90 4.36 0.5 5.80 1.75 4.65 4.40 4.70 4.75 4.34 0.75 5.70 3.05 4.80 4.40 4.95 5.05 4.66 1.0 5.65 1.40 4.75 4.40 4.65 4.50 4.22 GLM-Z1-9B 0.0 5.70 2.10 4.30 4.20 5.10 4.50 4.32 0.25 5.70 2.10 4.30 4.20 5.15 4.50 4.33 0.5 5.90 2.20 4.40 4.25 5.10 4.50 4.39 0.75 5.90 2.00 4.50 3.90 5.30 4.65 4.38 1.0 5.90 2.20 4.60 4.00 5.15 4.55 4.40 QwQ-32B 0.0 5.90 1.95 4.90 4.70 5.40 4.90 4.62 0.25 5.90 2.55 5.00 4.80 5.15 4.60 4.67 0.5 5.95 2.65 4.15 4.85 5.45 4.80 4.64 0.75 5.75 1.85 4.80 4.70 5.80 4.80 4.62 1.0 5.90 1.40 4.55 4.85 5.35 4.75 4.47 R1-Distill-70B 0.0 5.30 3.10 4.20 4.40 4.60 4.20 4.30 0.25 5.40 3.00 4.40 4.80 4.70 4.30 4.43 0.5 5.30 3.00 4.40 4.45 4.60 4.30 4.34 0.75 5.10 3.05 4.35 4.65 4.55 4.40 4.35 1.0 5.20 3.15 4.55 4.70 4.60 4.30 4.42 R1-Distill-8B 0.0 4.50 3.30 4.10 3.45 3.35 3.70 3.73 0.25 3.70 3.25 4.40 3.50 3.45 3.95 3.71 0.5 3.95 3.10 4.20 3.65 3.30 3.90 3.68 0.75 3.80 3.40 3.90 3.50 3.65 3.85 3.68 1.0 4.10 3.35 4.20 3.50 3.85 3.95 3.83 Qwen3-235B-R 0.0 5.60 2.70 4.40 4.75 4.95 4.40 4.47 0.25 5.75 2.80 4.90 4.35 4.85 4.45 4.52 0.5 5.50 2.70 4.60 4.65 4.80 4.60 4.47 0.75 5.50 2.85 4.70 4.30 5.00 4.50 4.47 1.0 5.55 3.00 4.90 4.75 5.15 4.40 4.62 Non-Reasoning Models Baichuan2-7B 0.0 4.00 2.60 3.20 3.05 3.20 3.65 3.28 0.25 3.45 2.40 3.25 3.60 3.40 3.70 3.30 0.5 3.40 2.95 3.00 3.60 3.90 4.00 3.48 0.75 2.25 2.60 3.15 3.60 3.45 3.30 3.06 1.0 3.90 2.55 4.00 3.45 3.05 4.05 3.50 DeepSeek-V3 0.0 5.00 3.30 4.30 3.95 4.10 4.00 4.11 0.25 4.90 3.30 4.30 3.75 4.20 4.00 4.08 0.5 5.10 3.30 4.00 3.80 4.10 3.90 4.03 0.75 5.20 3.30 4.00 3.80 4.10 3.95 4.06 1.0 5.10 3.30 4.00 3.80 4.10 3.90 4.03 GPT-4.1 0.0 5.60 3.30 4.40 5.40 4.80 4.20 4.62 0.25 5.65 3.30 4.40 5.40 4.75 4.20 4.62 0.5 5.70 3.30 4.40 5.40 4.80 4.20 4.63 0.75 5.60 3.30 4.40 5.40 4.80 4.20 4.62 1.0 5.60 3.30 4.45 5.40 4.90 4.20 4.64 Qwen2.5-32B 0.0 3.90 3.10 3.35 3.20 3.20 3.00 3.29 0.25 3.90 3.10 3.00 3.20 3.20 3.00 3.23 0.5 3.90 3.10 3.05 3.20 3.20 3.00 3.24 0.75 3.90 3.05 3.00 3.20 3.20 3.00 3.23 1.0 3.90 3.10 3.30 3.20 3.20 3.00 3.28 Qwen2.5-72B 0.0 4.95 2.70 3.90 3.70 4.20 4.00 3.91 0.25 4.80 2.70 3.90 3.70 4.20 4.00 3.88 0.5 4.75 2.70 3.90 3.70 4.20 4.00 3.88 0.75 4.80 2.70 3.85 3.75 4.20 3.95 3.88 1.0 4.80 2.75 3.90 3.70 4.20 4.00 3.89 Qwen2.5-7B 0.0 4.20 2.20 3.60 3.90 4.30 3.90 3.68 0.25 4.30 2.20 3.60 3.90 4.20 3.80 3.67 0.5 4.30 2.30 3.60 3.90 4.20 3.75 3.67 0.75 4.40 2.15 3.50 4.00 4.15 3.85 3.67 1.0 4.20 2.25 3.55 3.90 4.15 3.75 3.63 Qwen3-235B-C 0.0 5.20 3.05 4.80 4.70 4.90 4.40 4.51 0.25 5.30 3.10 4.90 4.70 4.85 4.40 4.54 0.5 5.25 3.20 4.90 4.75 5.00 4.40 4.58 0.75 5.25 3.15 4.90 4.70 4.90 4.50 4.57 1.0 5.30 3.20 4.70 4.60 5.00 4.30 4.52
Table 14: CPST-E Scores of Reasoning and Non-Reasoning LLMs across Temperatures